OpenStudy (timaashorty):

In general, matrix multiplication is not commutative. Show by counterexample that this is true. Then give an example of when matrix multiplication is commutative, Explain

OpenStudy (timaashorty):

@satellite73

OpenStudy (anonymous):

pick any matrix make is simple, like a two by two matrix multiply one way and then the other and see that you don't get the same thing

OpenStudy (timaashorty):

If I don't get the same thing it makes it not commutative or commutative?

OpenStudy (anonymous):

\[ \left( \begin{array}{cc} 1 & 2 \\ 3 & 4 \\ \end{array} \right)\times\left( \begin{array}{cc} 3 & 2 \\ 1 & 4 \\ \end{array} \right) \] for example try that one

OpenStudy (anonymous):

then try \[\left( \begin{array}{cc} 3 & 2 \\ 1 & 4 \\ \end{array} \right)\left( \begin{array}{cc} 1 & 2 \\ 3 & 4 \\ \end{array} \right)\]

OpenStudy (anonymous):

the result will not be the same that shows that \(A\times B\neq B\times A\) which shows they are NOT commutative

OpenStudy (timaashorty):

= [5 10] [13 22] =[9 14] [13 18] They're different

OpenStudy (timaashorty):

OHH!

OpenStudy (timaashorty):

But if they're still the same answer if multiplied both ways means they're commuative?

OpenStudy (anonymous):

yeah, that is what it means commutative means for all \(A, B\) \(AB=BA\) not true for matrices

OpenStudy (anonymous):

\[ \left( \begin{array}{cc} 1 & 2 \\ 2 & 2 \\ \end{array} \right).\left( \begin{array}{cc} 0 & 1 \\ 1 & 1 \\ \end{array} \right)=\left( \begin{array}{cc} 2 & 3 \\ 2 & 4 \\ \end{array} \right)\\ \left( \begin{array}{cc} 0 & 1 \\ 1 & 1 \\ \end{array} \right).\left( \begin{array}{cc} 1 & 2 \\ 2 & 2 \\ \end{array} \right)=\left( \begin{array}{cc} 2 & 2 \\ 3 & 4 \\ \end{array} \right) \]

OpenStudy (timaashorty):

Okay I see ^.^ Thank you soo much! You good helpers only come out at night ;P Now I just said need to find a multplication matrix that is commutative

OpenStudy (anonymous):

hint use the identity matrix it commutes with everything

OpenStudy (anonymous):

Take two diagonal matrices

OpenStudy (anonymous):

or what @eliassaab said

OpenStudy (timaashorty):

Is it okay if I take to same matrix and put them as an example? like this [1 2] [1 2] [1 4] [1 4]

OpenStudy (timaashorty):

I can't find one @eliassaab

OpenStudy (timaashorty):

Found one: [1 0] [3 0] [3 0] [0 2] [0 4] = [0 8] [3 0] [1 0] [3 0] [0 4] [0 2] = [0 8]